OpenStudy (cail):

(-5.1)^0 a. 1 << b. 0 c. -5.1 d. -1 3g^-2 b^2 a. 3b^2/g^2 b. 3g^2 b^-2 <<< c. 3gb^-4 d. b^2/3g^2 3/g^-2 h^3 a. 3/g^2 h^3 << b. 3g^2/h^3 c. 6g/h^3 d. 3/gh

OpenStudy (cail):

@mathstudent55 @RadEn

OpenStudy (cail):

@thomaster @math_genius123 @mathcrazy24

OpenStudy (raden):

only the first one is correct

OpenStudy (cail):

i changed the answer hold on

OpenStudy (cail):

two to a three to b

OpenStudy (raden):

yeah thats right :)

OpenStudy (cail):

thankks:)

OpenStudy (cail):

can you help with 2 more?

OpenStudy (raden):

i'll try

OpenStudy (cail):

What is the value of 2x^-2y^-2 for x=3 and y=-2 a. 1/18 b. 72 c. 2(-6)^-4 ******* d. 1/648 heres the first one

OpenStudy (cail):

What is the value of 1/2^-2x^-3y^5 a. 16 b. -4 c. -1/32 d. -16**

OpenStudy (raden):

hmmm... 2x^-2y^-2 for x=3 and y=-2 2x^-2y^-2 = 2/(x^2y^2) = 2/(xy)^2 = 2/(3*-2)^2 = 2/(-6)^2 = 2/36=1/18

OpenStudy (cail):

Oh so a.. my bad

OpenStudy (raden):

can you give the parantheses for the 2nd one, to make sure whats the original equation (or you can draw it)

OpenStudy (cail):

What do you mean?

OpenStudy (raden):

is it like this ? :